OpenStudy (anonymous):

Solve tan(sin^-1 1/2). Someone help please!

OpenStudy (anonymous):

do you knw the value to \(\large sin^{-1}(\frac{1}{2})= \) ???

OpenStudy (anonymous):

Is it 1/2 due to the inverse property??

OpenStudy (anonymous):

to find out, ask yourself this equivalent question: \(\large sin (x)=\frac{1}{2} \)

OpenStudy (anonymous):

sin(what angle) = 1/2 ??? are you familiar with the unit circle?

OpenStudy (anonymous):

page 3: http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf

OpenStudy (fwizbang):

draw a triangle with an angle whose sin =1/2. Use it to fiind the tangent.

OpenStudy (anonymous):

yes, some what-so 1/2=pi/6

OpenStudy (anonymous):

so now can you answer the question: \(\large tan(\frac{\pi}{6})= \) ???

OpenStudy (anonymous):

that will be equals 1/square root of 3

OpenStudy (anonymous):

yes... but you can also simplify that....:)

OpenStudy (anonymous):

yep just realized that lol...it will be square root of 3 over 3

OpenStudy (anonymous):

good job...:)

OpenStudy (anonymous):

=]

OpenStudy (anonymous):

thanks